OpenStudy (anonymous):

A tree farm initially has 5000 trees. Each year 10% of the trees are harvested and 450 seedlings are planted. Write a recursive rule for the number of trees on the tree farm at the beginning of the nth year.

OpenStudy (anonymous):

y(0)=5000 y(1)=y(0)-.1(y(0)+450 y(1)=.9y(0)+450 y(n)=.9y(n-1)+450